OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____. A. 3x = 2 B. 3x = 5 C. 5x = 5

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.

OpenStudy (amistre64):

which part of this is confusing you; the multiplying, or the adding?

OpenStudy (anonymous):

Multiplying... @amistre64

OpenStudy (amistre64):

lets take equation 1 2x + y = 3 and multiply it all by 2 2(2x + y = 3) 2(2x) + 2(y) = 2(3)

OpenStudy (anonymous):

Ok, i hate these types of problems cause i never understand them...

OpenStudy (amistre64):

do you understand now what multiply by 2 means?

OpenStudy (anonymous):

Yes now i do

OpenStudy (amistre64):

good ... and i assume you have the adding down pat :)